# Controllable Video Generation with Provable Disentanglement

### v2

@@ 5 Experiments
In this section, we perform extensive experimental validation to comprehensively evaluate CoVoGAN. We design both qualitative and quantitative experiments to systematically assess the generation quality and controllability of our proposed model. Finally, we perform comprehensive ablation studies to demonstrate the effectiveness of our proposed modules and analyze their individual contributions.
@@ 5.1 Experimental setup
Datasets. We evaluate our model on three different real-world datasets: FaceForensics rossler2018faceforensics , SkyTimelapse xiong2018learning , and RealEstate zhou2018stereo . These datasets respectively capture dynamic variations in facial expressions, sky transitions, and complex camera movements, providing comprehensive evaluation scenarios for our method. FaceForensics and SkyTimelapse are widely adopted benchmarks in video generation research, particularly for facial manipulation and natural scene modeling. The RealEstate dataset contains video sequences featuring complex camera movements within static scenes. The camera motions in RealEstate are more intuitive and easily interpretable by human observers, including forward/backward movements, lateral shifts, etc. We particularly leverage this dataset to effectively demonstrate the precise controllability of CoVoGAN in handling explicit camera motion parameters. All datasets consist of videos with a resolution of 256 × 256 pixels, and we employ standard train-test splits for fair evaluation. Detailed statistics and additional information about the datasets are provided in Appendix C.
Evaluation metrics. To comprehensively evaluate the performance of CoVoGAN, we employ both quantitative and qualitative assessment metrics. For quantitative evaluation, we adopt the Fréchet Video Distance (FVD) unterthiner2018towards , a widely-used metric for assessing video generation quality. FVD computes the Wasserstein-2 distance between feature distributions of real and generated videos in a learned feature space, providing a comprehensive measure of visual quality and temporal coherence. We report FVD scores at two different temporal scales: $\text{FVD}_{8}$ and $\text{FVD}_{16}$, where the subscript denotes the number of frames in a video. To better assess the disentanglement capability of our method, we compare SAP kumar2017variational and modularity ridgeway2018learning . Additionally, we measure the Mean Correlation Coefficient (MCC), a standard metric in the Independent Component Analysis (ICA) literature for continuous variables. For qualitative evaluation, we present randomly sampled video sequences generated by our model.
@@ 5.2 Video quality
We conduct comparisons between our proposed CoVoGAN and four GAN-based models: MoCoGAN-HD, DIGAN, StyleGAN-V, and MoStGAN-V, across three benchmark datasets: FaceForensics, SkyTimelapse, and RealEstate. We also compare three diffusion-based models: VDM, LVDM he2022latent and Latte ma2024latte . For MoCoGAN-HD, we utilize the image generator pretrained by the original authors and train the remaining components (including the motion generator and discriminator). Due to the unavailability of pretrained models for the RealEstate dataset, we exclude MoCoGAN-HD from the comparisons on this specific dataset. For Latte, we directly use the released checkpoints. For the remaining baselines, we train the models from scratch using their official implementations to ensure comparability. The quantitative results in Table 1 show that CoVoGAN consistently achieves top performance across all datasets, despite diffusion-based models occasionally exhibiting higher visual fidelity. Additional randomly sampled generations are provided in the Appendix E.2.
Figure 4: Disentangled control over identity and motion in (a) FaceForensics, (b) RealEstate.
Table 1: $\text{FVD}_{8}\downarrow$ and $\text{FVD}_{16}\downarrow$ scores across the FaceForensics, SkyTimelapse, and RealEstate datasets.
Method FaceForensics $\text{256}^{2}$ SkyTimelapse $\text{256}^{2}$ RealEstate $\text{256}^{2}$
$\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{8}$ $\text{FVD}_{16}$
MoCoGAN-HD 140.05 185.51 1214.13 1721.89 - -
@@
Latte 45.49 49.02 40.21 41.84 - -
CoVoGAN (ours) 43.75 48.80 35.58 46.51 154.88 174.87
@@ 5.3 Controllability
For CoVoGAN, we modify one dimension of the style dynamics ${\mathbf{z}}_{t}^{s}$. For StyleGAN-V and MoStGAN-V, we manipulate one dimension of their (latent) motion code. We apply equivalent modifications to the corresponding latent dimensions in each baseline model. To validate the consistency of our controllability analysis, we randomly sample three distinct video sequences and apply identical
Figure 5: Controllability visualization results on the FaceForensics dataset. Two different concepts of motion are modified to show the component-wise level disentanglement. The first line is a randomly sampled video, and the following lines apply control of eye blinking and head shaking. Each subplot displays a small green arrow indicating the measured distance between the facial contour and the image boundary.
modifications to their respective latent representations. Other baselines are not compared since there are no specific variables with semantic information.
@@
The results show that our proposed CoVoGAN model learns a disentangled representation that effectively separates style dynamics from content elements. (1) This disentanglement enables independent manipulation of motion characteristics while preserving content consistency. (2) A key advantage of this approach is that identical modifications to the style latent space consistently produce similar motion patterns across different content identities.
Baseline models achieve only partial disentanglement, exhibiting two major limitations: (1) visual distortions of the modified videos and (2) inconsistent or misaligned motion patterns when applied to different identities.
@@
Component-wise disentanglement. We further illustrate the precise control over individual motion components. This capability is enabled by the component-wise identifiability of our model, which ensures each latent dimension corresponds to a specific and interpretable motion attribute, e.g., eye blinking or head shaking.
@@
Our experimental procedure begins with randomly sampling two distinct video sequences, as illustrated in the first row of Figure 5. We then selectively modify the latent dimension corresponding to eye blinking dynamics in the second line. Subsequently, we modify a second latent dimension controlling head shaking motion while maintaining the previously adjusted eye blinking pattern in the last line. The results show naturalistic head movements from left to right, synchronized with the preserved eye blinking, illustrating our model’s capability for independent yet coordinated control of multiple motion components.
Table 2: Comparison of disentanglement metrics across different models. Higher is better.
Metrics CoVoGAN StyleGAN-V MoStGAN-V
MCC (%) $\uparrow$ 33.78 29.00 27.95
SAP (%) $\uparrow$ 8.48 4.25 5.90
Modularity (%) $\uparrow$ 17.37 7.66 13.48
@@
Disentanglement metrics. To evaluate disentanglement performance, we first extract semantic annotations from the videos. For this purpose, we conduct experiments on the FaceForensics dataset, which facilitates the extraction of meaningful facial attributes. We utilize a pretrained model from Dlib to extract facial landmarks and compute semantic annotations, including eye size, mouth size, head position, and head angle. We only compare our method with two models that explicitly incorporate semantic representation layers, which we use to compute the metrics. Diffusion-based models without semantic representations are not included in this comparison. For our method, we compute the metrics using the dynamic latent variables $z^{s}_{t}$. As shown in Table 2, our method achieves the best performance across all disentanglement metrics.
@@ 5.4 Ablation study
We conduct an ablation study to evaluate the contributions of the proposed components within the Temporal Transition Module, as shown in Table 3.
Replacing the GRU with a standard RNN leads to a noticeable performance drop. This is primarily due to the loss of sparsity provided by the GRU’s gating mechanism, which helps isolate time-delayed effects. Without this mechanism, the model faces a larger search space, making it harder to learn an effective transition function.
Substituting the component-wise flow with a fully connected MLP results in an even more significant degradation. This decline can be attributed to two factors: (1) the mutual independence between style dynamics can no longer be maintained, and (2) capturing sufficient changes becomes more difficult.
Table 3: Ablation study results for different configurations of GRU and component-wise flow on the FaceForensics dataset.
@@ 6 Conclusion
In this paper, we proposed a Temporal Transition Module and implemented it in a GAN to achieve CoVoGAN. By leveraging the principles of minimal and sufficient changes, we successfully disentangled (1) the motion and content of a video, and (2) different concepts within the motion. We established an identifiability guarantee for both block-wise and component-wise disentanglement. Our proposed CoVoGAN model demonstrates high generative quality and controllability. We validated the performance on various datasets and conducted ablation experiments to further confirm the effectiveness of our model. Overall, our work provides a principled and practical framework for disentangled video generation, and opens new directions for fine-grained, interpretable, and controllable visual synthesis. Nonetheless, our study has certain limitations. It primarily centers on theoretical contributions and their empirical validation, while the integration with higher-fidelity generative architectures and state-of-the-art text-to-video synthesis frameworks remains an open direction for future research.
@@ A.1 Proof
Monoblock Identification. Following [10, 9], when assumptions B1, B2, B3 satisfied, the blockwise identifiability of $[{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]$ is assured, according to Theorem 3.2(Monoblock identifiability) in [6]. In short ,there exists a invertible function $g$ such that $[{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$, where []delimited-[][$\cdot$][ ⋅ ] denotes concatenation.
@@
$[{\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h([\hat{\mathbf{z}}_{t}^{s},\hat{\mathbf{z}}^{c}]),$ (12)
where [.][.][ . ] denotes the concatenation operation. The corresponding Jacobian matrix can be formulated as
@@
[…]
@@
Next, using the mutual independence of $\hat{{\mathbf{z}}}_{t}^{s}|\hat{{\mathbf{z}}}_{t-1}^{s}$ in assumption C3, we have $\frac{\partial^{2}\hat{\mathbb{\eta}}_{t}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=0$ according to the connection between conditional independence and cross derivatives [24]. Thus we have
@@ A.2 Discussion of Assumptions
Assumption B1 requires a smooth distribution, which is usually hold in the real world. Secondly, Assumption B2 imposes a minimal requirement on the number of variables. The linear operator $L_{b|a}$ ensures that there is sufficient variation in the density of $b$ for different values of $a$, thereby guaranteeing injectivity. In a video, ${\mathbf{x}}_{t}$ is of much higher dimensionality compared to the latent variables. As a result, the injectivity assumption is easily satisfied. In practice, following the principle of minimal changes, if a model with fewer latent variables can successfully achieve observational equivalence, it is more likely to learn the true distribution. Assumption B3 requires the distribution of ${\mathbf{x}}_{t}$ changes when the value of latent variables changes. This assumption is much weaker compared to the widely used invertibility assumption adopted by previous works, such as [52]. The aforementioned assumptions concern the underlying data-generating process. The aforementioned assumptions pertain to the underlying data-generating process, and are often easily satisfied in real-world video scenarios. Given a model that can explicitly disentangle motion and identity while maintaining high generative fidelity, block-wise disentanglement becomes a natural and attainable property under these conditions. This also explains why models such as MoCoGAN [37] are able to effectively separate motion and identity.
Assumption C1 further requires the underlying distribution to be smooth, while Assumptions C2 and C3 impose mild conditions on the variability of latent factors. These assumptions are not overly restrictive and are often satisfied in practice. Notably, even when the sufficiency conditions are satisfied only by a subset of latent variables, or when mutual independence is partially violated, identifiability can still be achieved at the subspace level, as demonstrated by [17] in Section A.1. To further attain disentanglement at the component-wise level, the architectural design of our proposed Temporal Transition Module (TTM) is essential.

### v3

@@ 5.1 Experimental setup
Datasets. We evaluate our model on four different real-world datasets: FaceForensics (Rössler et al., 2018), SkyTimelapse (Xiong et al., 2018), RealEstate (Zhou et al., 2018) and CelebV-HQ (Zhu et al., 2022). The first three datasets contain videos with a resolution of 256 × 256 pixels, while the last dataset consists of videos at 512 × 512 resolution. We employ standard train-test splits for fair evaluation. Detailed information about the datasets are provided in Appendix C.
Figure 4: Controllability in the latent space across datasets and methods. Each method is evaluated with three samples by varying a single latent dimension. Only CoVoGAN exhibits consistent control across identities: (a) head pose adjustment on FaceForensics, (b) camera translation on RealEstate.
Evaluation metrics. To comprehensively evaluate the performance of CoVoGAN, we employ both quantitative and qualitative assessment metrics. For quantitative evaluation, we adopt the Fréchet Video Distance (FVD) (Unterthiner et al., 2018), a widely-used metric for assessing video generation quality. We report FVD scores at two different temporal scales: $\text{FVD}_{8}$ and $\text{FVD}_{16}$, where the subscript denotes the number of frames in a video. To better assess the disentanglement capability of our method, we compare the widely used disentanglement metrics SAP (Kumar et al., 2017) and modularity (Ridgeway & Mozer, 2018). Additionally, we measure the Mean Correlation Coefficient (MCC), a standard metric for disentanglement. Details of the metrics can be found in Appendix D.
@@ 5.2 Video quality
Table 1: $\text{FVD}_{8}\downarrow$ and $\text{FVD}_{16}\downarrow$ results across different datasets.
Method FaceForensics SkyTimelapse RealEstate CelebV-HQ
$\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{8}$ $\text{FVD}_{16}$ $\text{FVD}_{16}$
MoCoGAN-HD 140.05 185.51 1214.13 1721.89 412.50
@@
Latte 45.49 49.02 40.21 41.84
CoVoGAN 43.75 48.80 35.58 46.51 154.88 174.87 97.16
We consider four GAN-based models: MoCoGAN-HD, DIGAN, StyleGAN-V, and MoStGAN-V. We also compare three diffusion-based models: VDM (Ho et al., 2022), LVDM (He et al., 2022) and Latte (Ma et al., 2024). For MoCoGAN-HD, we freeze the image generator pretrained by the original authors and train the motion generator and discriminator. We exclude MoCoGAN-HD from the comparisons on the specific datasets without pretrained image generator. For Latte, we directly use the released checkpoints. For the remaining baselines, we train the models from scratch using their official implementations to ensure comparability. The quantitative results in Table 1 show that CoVoGAN consistently achieves top performance across all datasets, despite diffusion-based models occasionally exhibiting higher visual fidelity. More visualization results are provided in the Appendix F.5.
@@ 5.3 Controllability
Figure 5: Controllability visualization results on the FaceForensics dataset. Two distinct motion concepts are manipulated to illustrate component-wise disentanglement. Corresponding videos are provided in the supplementary materials for better visualization.
@@
For CoVoGAN, we modify one dimension of the style dynamics ${\mathbf{z}}_{t}^{s}$. For StyleGAN-V and MoStGAN-V, we manipulate one dimension of their (latent) motion code. We apply equivalent modifications to the corresponding latent dimensions in each baseline model. To validate the consistency of our controllability analysis, we randomly sample three distinct video sequences and apply identical modifications to their respective latent representations. Other baselines are not compared since there are no specific variables with semantic information.
@@
The results show that our proposed CoVoGAN model learns a disentangled representation that effectively separates style dynamics from content elements. (1) This disentanglement enables independent manipulation of motion characteristics while preserving content consistency. (2) A key advantage of this approach is that identical modifications to the style latent space consistently produce similar motion patterns across different content identities. Baseline models achieve only partial disentanglement, exhibiting two major limitations: (1) visual distortions of the modified videos and (2) inconsistent or misaligned motion patterns when applied to different identities.
@@
Component-wise disentanglement. We also show the precise control over individual motion components. This capability is enabled by the component-wise identifiability of our model,
Table 2: Disentanglement Comparison across different models.
Metrics StyleGAN-V MoStGAN-V LVDM Latte CoVoGAN
MCC (%) $\uparrow$ 29.00 27.95 21.60 20.87 33.78
SAP (%) $\uparrow$ 4.25 5.90 0.72 0.75 8.48
Modularity (%) $\uparrow$ 7.66 13.48 7.25 7.44 17.37
@@
which ensures each latent dimension corresponds to a specific and interpretable motion attribute, e.g., eye blinking or head shaking. Our experimental procedure begins with randomly sampling two distinct video sequences, as illustrated in the first row of Figure 5. We then selectively modify the latent dimension corresponding to eye blinking dynamics in the second line. Subsequently, we modify a second latent dimension controlling head shaking motion while maintaining the previously adjusted eye blinking pattern in the last line. The results show naturalistic head movements from left to right, synchronized with the preserved eye blinking, illustrating our model’s capability for independent yet coordinated control of multiple motion components.
@@
Disentanglement metrics. We first extract semantic annotations from the videos for disentanglement evaluation. For this purpose, we conduct experiments on the FaceForensics dataset, which facilitates the extraction of meaningful facial attributes. We utilize a pretrained model from Dlib to extract facial landmarks and compute semantic annotations, including eye size, mouth size, head position, and head angle. We only compare our method with two models that explicitly incorporate semantic representation layers, which we use to compute the metrics. Since diffusion-based models do not provide a compact latent representation by design, we adopt the following procedure: we first randomly sample videos, then extract frame-wise representations from the high-dimensional latent space. These representations are reduced to 128 dimensions using PCA, and the disentanglement metrics are computed in this reduced space. For our method, we compute the metrics using the dynamic latent variables $z^{s}_{t}$. As shown in Table 2, our method achieves the best performance.
@@ 5.4 Ablation study
Table 3: Ablation studies on different GRU configurations and component-wise flow on the FaceForensics dataset.
@@
We conduct an ablation study to evaluate the contributions of the proposed components within the Temporal Transition Module, as shown in Table 3. Replacing the GRU with a standard RNN leads to a noticeable performance drop. This is primarily due to the loss of sparsity provided by the GRU’s gating mechanism, which helps isolate time-delayed effects. Without this mechanism, the model faces a larger search space, making it harder to learn an effective transition function. Substituting the component-wise flow with a fully connected MLP results in an even more significant degradation. This decline can be attributed to two factors: (1) the mutual independence between style dynamics can no longer be maintained, and (2) capturing sufficient changes becomes more difficult.
@@ 6 Conclusion
In this paper, we proposed a Temporal Transition Module and implemented it in a GAN to achieve CoVoGAN. By leveraging the principles of minimal and sufficient changes, we successfully disentangled (1) the motion and content, and (2) different concepts within the motion. We established an identifiability guarantee for both block-wise and component-wise disentanglement. Our proposed CoVoGAN model demonstrates high generative quality and controllability. We validated the performance on various datasets and conducted ablation experiments to further confirm the effectiveness of our model. Overall, our work provides a principled and practical framework for disentangled video generation and opens new directions for fine-grained, interpretable, and controllable visual synthesis. Limitations and Future works: It focuses primarily on theoretical contributions and their empirical validation, while the integration with higher-fidelity generative architectures and application to open-domain scenarios remains an important direction for future work.
Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. A complete description of the datasets and preprocessing steps is included in Appendix C. The implementation details, including model architectures, training hyperparameters, and optimization settings, are provided in Appendix E. Extended results are reported in Appendix F. In addition, we release the full source code and configuration files in the supplementary materials.
@@ A.1 Proof
Monoblock Identification. Following (Hu & Shum, 2012; Hu & Schennach, 2008b), when assumptions B1, B2, B3 satisfied, the blockwise identifiability of ${\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]$ is assured, according to Theorem 3.2 (Monoblock identifiability) in (Fu et al., 2025). In short ,there exists a invertible function $g$ such that ${\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c})$, where $\cdot$] denotes concatenation.
@@
${\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}]=h([\hat{\mathbf{z}}_{t}^{s},\hat{\mathbf{z}}^{c}]),$ (12)
where $[.]$ denotes the concatenation operation. The corresponding Jacobian matrix can be formulated as
@@
$\begin{aligned} &&p({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}|{\mathbf{x}}_{t-1})&=p(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}|{\mathbf{x}}_{t-1})/|H_{t}|\\ &\Rightarrow&p({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}|g({\mathbf{z}}_{t-1}^{s},{\mathbf{z}}^{c}))&=p(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}|\hat{g}(\hat{{\mathbf{z}}}_{t-1}^{s},\hat{{\mathbf{z}}}^{c}))/|H_{t}|\\ &\Rightarrow&p({\mathbf{z}}_{t}^{s},{\mathbf{z}}^{c}|{\mathbf{z}}_{t-1}^{s},{\mathbf{z}}^{c})&=p(\hat{{\mathbf{z}}}_{t}^{s},\hat{{\mathbf{z}}}^{c}|\hat{{\mathbf{z}}}_{t-1}^{s},\hat{{\mathbf{z}}}^{c})/|H_{t}|\\ &\Rightarrow&p({\mathbf{z}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s},{\mathbf{z}}^{c})&=p(\hat{{\mathbf{z}}}_{t}^{s}|\hat{{\mathbf{z}}}_{t-1}^{s},\hat{{\mathbf{z}}}^{c})/|H_{t}|\\ &\Rightarrow&p({\mathbf{z}}_{t}^{s}|{\mathbf{z}}_{t-1}^{s})&=p(\hat{{\mathbf{z}}}_{t}^{s}|\hat{{\mathbf{z}}}_{t-1}^{s})/|H_{t}|\\ \end{aligned}.$ (15)
@@
Next, using the mutual independence of $\hat{{\mathbf{z}}}_{t}^{s}|\hat{{\mathbf{z}}}_{t-1}^{s}$ in assumption C3, we have $\frac{\partial^{2}\hat{\mathbb{\eta}}_{t}}{\partial\hat{z}_{t,i}^{s}\partial\hat{z}_{t,j}^{s}}=0$ according to the connection between conditional independence and cross derivatives (Lin, 1997). Thus we have
@@ A.2 Discussion of Assumptions
Assumption B1 requires a smooth distribution, which is usually hold in the real world. Secondly, Assumption B2 imposes a minimal requirement on the number of variables. The linear operator $L_{b|a}$ ensures that there is sufficient variation in the density of $b$ for different values of $a$, thereby guaranteeing injectivity. In a video, ${\mathbf{x}}_{t}$ is of much higher dimensionality compared to the latent variables. As a result, the injectivity assumption is easily satisfied. In practice, following the principle of minimal changes, if a model with fewer latent variables can successfully achieve observational equivalence, it is more likely to learn the true distribution. Assumption B3 requires the distribution of ${\mathbf{x}}_{t}$ changes when the value of latent variables changes. This assumption is much weaker compared to the widely used invertibility assumption adopted by previous works, such as (Yao et al., 2022). The aforementioned assumptions concern the underlying data-generating process. The aforementioned assumptions pertain to the underlying data-generating process, and are often easily satisfied in real-world video scenarios. Given a model that can explicitly disentangle motion and identity while maintaining high generative fidelity, block-wise disentanglement becomes a natural and attainable property under these conditions. This also explains why models such as MoCoGAN (Tulyakov et al., 2018) are able to effectively separate motion and identity.
Assumption C1 further requires the underlying distribution to be smooth, while Assumptions C2 and C3 impose mild conditions on the variability of latent factors. These assumptions are not overly restrictive and are often satisfied in practice. Notably, even when the sufficiency conditions are satisfied only by a subset of latent variables, or when mutual independence is partially violated, identifiability can still be achieved at the subspace level, as demonstrated by (Kong et al., 2023) in Section A.1. To further attain disentanglement at the component-wise level, the architectural design of our proposed Temporal Transition Module (TTM) is essential.
A.3 Examples of Injective Linear Operators
The assumption that a linear operator is injective is commonly used in nonparametric identification (Hu & Schennach, 2008b; Carroll et al., 2010; Hu & Shum, 2012). Intuitively, this means that distinct input distributions of a linear operator correspond to distinct output distributions. To clarify this assumption, we provide several examples illustrating the mapping from $p_{{\mathbf{a}}}\rightarrow p_{{\mathbf{b}}}$, where $\mathbf{a}$} and $\mathbf{b}$} are random variables.
Example 1 (Inverse Transformation).
$b=g(a)$, where $g$ is an invertible function.
Example 2 (Additive Transformation).
$b=a+\epsilon$, where the distribution $p(\epsilon)$ must not vanish entirely under the Fourier transform (Theorem 2.1 in (Mattner, 1993)).
Example 3.
$b=g(a)+\epsilon$, requiring the same conditions as in Examples 1 and 2.
Example 4 (Post-linear Transformation).
$b=g_{1}(g_{2}(a)+\epsilon)$, a post-nonlinear model with invertible nonlinear functions $g_{1}$ and $g_{2}$, combining the assumptions in Examples 13.
Example 5 (Nonlinear Transformation with Exponential Family).
$b=g(a,\epsilon)$, where the joint distribution $p(a,b)$ belongs to an exponential family.
Example 6 (General Nonlinear Transformation).
$b=g(a,\epsilon)$, representing a general nonlinear mapping. Certain deviations from the nonlinear additive model (Example 3), such as polynomial perturbations, can still be tractable.